OpenStudy (anonymous):

If theta is any real number, is it possible for sin theta = 5/4 ? Why or why not?

jimthompson5910 (jim_thompson5910):

hints 5/4 = 1.25 \[\Large -1 \le \sin(\theta) \le 1\]

OpenStudy (anonymous):

so you all you did was divide the fraction? @jim_thompson5910

OpenStudy (phi):

Jim is saying that the sin of an angle is between -1 and 1 He also pointed out that 5/4 written as a decimal is 1.25, to make it more clear that 5/4 is bigger than 1 (and so outside the legal values for sin)

jimthompson5910 (jim_thompson5910):

yes I used a calculator to compute 5/4 = 1.25 you can also use long division

jimthompson5910 (jim_thompson5910):

the second thing basically says that sin(theta) is between -1 and 1 The smallest that sin(theta) can get is -1 The largest that sin(theta) can get is +1